Please fill the table for your shift with this information:

  • PROGRAM DONE: please write the type of measurements you have done, with reference to run number range and to logbook pages
  • BEAM INFO: note periods with no beam, report related problems
  • ASIDE ACTIVITIES: please add information on aside activities such as debugging… put references to logbook pages
  • OBSERVED PROBLEMS: add Meaningful information of problems encountered in your shift

Shift Date 24.11.2012 - Saturday
Shift Time 16-24
Shifter Names Richard, Guido
Program Done calo towers calibration-towers 8/9/10/14/15/16/20/21/22- S pms calibrated at 10%; C pms gain quite low for several towers with HV at 900
Beam Info no beam till 5 pm
Aside Activities neutron counters: some lemo cables replaced
Observed Problems DAQ problem; PC changed- restart around 7 pm

Shift Date 25.11.2012 - Sunday
Shift Time 0-8
Shifter Names Sehwook, Joao
Program Done Calibration for calorimeter towers: 1,2,3,4,5,11,17,23,29,28,27,26,25,19,13,7 - C and S channels were calibrated at 10%
Beam Info 20 GeV e-
Aside Activities None
Observed Problems S11 showed overflow and C11 has only pedestal (Run 9000)

Shift Date 25.11.2012 - Sunday
Shift Time 8-16
Shifter Names Alan, John
Program Done calibrated the blue perimeter; re-calibrated 4 C and 1 S tapered channel; calibrated n-counters 9,10,11,12 & 1,2,3,4 & 5,6,7,8.
Beam Info 20 GeV e-
Aside Activities  
Observed Problems cannot reach 8 n-counters on top with table.

Shift Date 25.11.2012 - Sunday
Shift Time 16-24
Shifter Names Andy, Gabri
Program Done Finished calibrating neutron counters. Connected the central tower to the TDC. Installed the interaction target. Started data taking with 60GeV pions.
Beam Info Stable beam.
Aside Activities Resolution changed the TDC setting from 0.8472 ns/cts to 0.14 ns/cts.
Observed Problems The DAQ stuck a few times because no spill signal.

Shift Date 26.11.2012 - Monday
Shift Time 0-8
Shifter Names Richard, Tim
Program Done  
Beam Info  
Aside Activities  
Observed Problems  

Shift Date 26.11.2012 - Monday
Shift Time 8-16
Shifter Names Evelin, Guido
Program Done Electron measurements started: taken a long run of 100kevts (set of small datasets) with 20 GeV ele at center (tower 15), nominal position
Beam Info Some problems with RF, stopped the runs during periods without beam
Aside Activities Around 2 p.m. stopped last run, and started operation to tilt New Dream for the electron energy scan
Observed Problems Impossible to set a pion beam of 40 GeV, so skipped pions measurements and started electron measurements

Shift Date 26.11.2012 - Monday
Shift Time 16-24
Shifter Names Gabri, Marco
Program Done Found T15 coordinates after calorimeter was tilted (logbook page 60). Started electron energy scan (logbook page 61)
Beam Info no beam for most of the shift
Aside Activities Prepared cables for the Copper module readout.
Observed Problems ---

Shift Date 27.11.2012 - Tuesday
Shift Time 00-08
Shifter Names Joao,Laura
Program Done energy scan 20GeV, 6 GeV (veto out), 40 GeV (running).
Beam Info no beam before 2.00
Aside Activities Connection of Cu module to HV power supply, Board 06. Channel C2 do not draw any current
Observed Problems Low voltage was off. BC voltage off.

Shift Date 28.11.2012-Wednesday
Shift Time 00-08
Shifter Names Andy, Sehwook
Program Done Cu(Al), Cu moudule calibration - Al1, Al2, Al3, Al4, Cu1, Cu2, Cu4. Energy scan with electron: 20 GeV(50k), 40 GeV(100k), 30 GeV(50k).
Beam Info stable beam
Aside Activities None
Observed Problems DAQ stopped data taking. Power-cycled VME crates.

Shift Date 28.11.2012-Wednsday
Shift Time 08-16
Shifter Names Guido,Tim
Program Done 6 Gev e- on copper module /// calo back at 0- 20 Gev e- on cu module to measure shower profile
Beam Info smooth running
Aside Activities vertical and angular titlt of the calo removed
Observed Problems none

Shift Date 29.11.2012 - Thursday
Shift Time 00-08
Shifter Names Laura, Gabri
Program Done 20 GeV pion run in the lead central tower (T15). Logbook page 71, Runs 9250 - 9267; Run 9268 20 GeV in T8
Beam Info 3:30 low current on target due to injection problem; then back to 25/e11; 4:25 - 5:30 no beam; then 25/e11
Aside Activities debugged channel T15S for timing measurements: wrong cable connected to the patch panel in CR; Changed back to NIM discriminator
Observed Problems At 7000 evt/spill DAQ frequently crashed. Rate reduced to 3000 evt/spill, ok

Shift Date 29.11.2012 - Thursday
Shift Time 16-24
Shifter Names  
Program Done  
Beam Info  
Aside Activities  
Observed Problems  

Shift Date 30.11.2012-Friday
Shift Time 00-08
Shifter Names Fabrizio, Alan
Program Done Pb and Cu(Al) modules, special runs with electrons
Beam Info 60, 80 and 100 GeV e-, stable
Aside Activities  
Observed Problems Gate problems with ancillary detectors, fixed through power cycling of crates and reloading FPGA software

Shift Date 30.11.2012-Friday
Shift Time 08-16
Shifter Names Tim, Laura
Program Done energy scan electron beam cu&pb modules ended (pg. 75 logbook). Calorimeter at (0,0). IT in. Crystal detectors in place.
Beam Info some instabilities
Aside Activities  
Observed Problems  

Shift Date 30.11.2012
Shift Time 16-24
Shifter Names Guido, Marco
Program Done calibration of the upper line of the neutron counters(13-20)- beam on tower 5- start 2M evts data taking-S5 saturation for deep shower -pag 77-79s
Beam Info 180 pai+ veto in 7-8K/spill
Aside Activities DRS short run
Observed Problems DAQ stop 3/4 times at 10K beam intensity

Shift Date 1.12.2012
Shift Time 0-8
Shifter Names Andy, John
Program Done Finished 2Mevts for lead module and completed 2Mevts for copper module (with saturation).
Beam Info 180 pi+
Aside Activities  
Observed Problems Several DAQ problems

Shift Date 1.12.2012
Shift Time 8-16
Shifter Names Tim, Alan
Program Done 2Mevt for copper and lead modules with adjusted voltages for CuAl2 and T15.
Beam Info 180 pi+
Aside Activities  
Observed Problems DAQ would stop, fixed with power cycle.

Shift Date 1.12.2012
Shift Time 16-24
Shifter Names Guido, Marco, Gabri
Program Done Calibration with lower HV in T15 and Al-Cu2 using 80 GeV e+ beam; finding crystal positions;
Beam Info stable beam
Aside Activities none
Observed Problems none

Shift Date 2.12.2012
Shift Time 00-08
Shifter Names Sehwook,Laura
Program Done Crystals measurement with different Pb depth at the front (electrons 80 Gev)- (logbook pg. 90-100
Beam Info some instabilities
Aside Activities no activity
Observed Problems Data were non correctly transferred and saved.Alessandro rebooted pcdream, pcdreamdaq2, and dreamix2. Voltage went suddenly off on DWC1.

Shift Date  
Shift Time  
Shifter Names  
Program Done  
Beam Info  
Aside Activities  
Observed Problems  
Edit | Attach | Watch | Print version | History: r24 < r23 < r22 < r21 < r20 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r24 - 2012-12-02 - LauraLaRotonda
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    DREAM All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback